OpenStudy (anonymous):

Which expression is equivalent to sin2x−sinx−2/sinx−2 ? sin x + 1 sin x − 1 sin2x -sin2x

OpenStudy (usukidoll):

I'm a bit confused... for the sin2x part is it really sin2x or (sinx)^2 -> sin^2x

OpenStudy (usukidoll):

\[\large \frac{\sin^2x-sinx-2}{sinx-2} \] like this?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

i get a decimal but there isn't decimals

OpenStudy (usukidoll):

yeah that would work... it's just like factoring only with trigs attached to it \[\large \frac{(sinx-2)(sinx+1)}{sinx-2}\] cancel like terms and see what happens

OpenStudy (anonymous):

sinx+1

OpenStudy (usukidoll):

:) yes

OpenStudy (anonymous):

oh wow i was so over thinking that lol thank you so much for your help